# Evaluating the Performance and Robustness of LLMs in Materials Science Q&A and Property Predictions

### v1

@@ Abstract
Large Language Models (LLMs) have the potential to revolutionize scientific research, yet their robustness and reliability in domain-specific applications remain insufficiently explored. This study conducts a comprehensive evaluation and robustness analysis of LLMs within the field of materials science, focusing on domain-specific question answering and materials property prediction. Three distinct datasets are used in this study: 1) a set of multiple-choice questions from undergraduate-level materials science courses, 2) a dataset including various steel compositions and yield strengths, and 3) a band gap dataset, containing textual descriptions of material crystal structures and band gap values. The performance of LLMs is assessed using various prompting strategies, including zero-shot chain-of-thought, expert prompting, and few-shot in-context learning. The robustness of these models is tested against various forms of ‘noise’, ranging from realistic disturbances to intentionally adversarial manipulations, to evaluate their resilience and reliability under real-world conditions. Additionally, the study uncovers unique phenomena of LLMs during predictive tasks, such as mode collapse behavior when the proximity of prompt examples is altered and performance enhancement from train/test mismatch. The findings aim to provide informed skepticism for the broad use of LLMs in materials science and to inspire advancements that enhance their robustness and reliability for practical applications.
Introduction
@@
Large Language Models (LLMs) represent a significant advancement in the field of artificial intelligence and have been rapidly adopted for application in various scientific disciplines [1]. With their ability to process and generate natural language, LLMs are potent tools for tasks like information retrieval, question and answering (Q&A), and property predictions [2, 3, 4]. Similar to traditional ML models, LLMs can require extensive data processing, large volumes of data, and massive compute resources to train [5]. Despite these limitations, pretrained LLMs can be adapted to new tasks with few-shot examples via in-context learning (ICL), making them both cost-effective and rapid to deploy [6, 7]. In the context of materials science, where data acquisition can often be costly and time-consuming, leveraging ICL enables LLMs to efficiently prototype and generate predictive insights even in low-data settings [8, 9, 10]. Recent work has demonstrated that LLMs are capable of domain-specific Q&A [11, 12, 13], materials property predictions [14, 15, 16], and information extraction from complex datasets [4, 17]. These studies demonstrate LLMs’ potential to serve as flexible and powerful analytical tools.
@@
However, the robustness of LLMs is crucial in real-world scenarios, but has not been sufficiently investigated in the literature. One aspect of the robustness of LLMs is their sensitivity to prompt changes either due to innocuous or adversarial reasons [18, 19]. Variations in how a query or instruction is formulated may cause a response to factually change [18]. As an example, 0.1 nm and 1 Angstrom are equivalent but switching them in a prompt could result in different LLM predictions for the same task. Alternatively, the response of the LLM can be deliberately altered through intentional misinformation or misleading inputs [19]. These attributes are not only theoretical concerns but are critical for the reliable usage of LLMs as they become integrated into the materials science research and development pipeline. Given that LLMs generate outputs with indifference to truth [20], thoroughly probing LLM prompt sensitivity would allow us to critically evaluate model performance in practical situations; providing informed skepticism for the broad use of LLMs in materials science.
@@
In this work, we conducted a holistic robustness analysis of commercial and open-source LLMs for materials science. We systematically investigated how textual perturbations like text ordering within a prompt, and common unintentional and intentional prompt variations influence the inference of LLMs. Three distinct datasets of domain-specific Q&A and materials property prediction were selected and first benchmarked via prompt engineering to identify baseline and optimal performance boundaries. We then investigated the impact of different types of textual perturbations, from realistic to adversarial perturbations, on the performance of various-sized LLMs in the materials science Q&A. We then used property predictions from the matbench_steels dataset to show that even without fine-tuning, pretrained LLMs demonstrated enhanced predictive ability through few-shot ICL when presented with similar examples to the prediction task. Conversely, when provided with dissimilar examples during few-shot ICL, mode collapse behavior was observed, where the model generated identical outputs despite varying inputs. Counterintuitively, for the fine-tuned LLM examined on band gap predictions, supposedly adversarial perturbations like shuffling or randomization enhanced the model’s predictive capability. Interestingly, this enables the fine-tuned model to function effectively even with significantly ablated prompts. This train/test mismatch behavior, absent in traditional ML models, highlights a potential direction for distilling LLM-based predictive models.
Results
@@
An overview of the framework is illustrated in Figure 1. For the performance evaluation and robustness analysis of materials science Q&A, we use the MSE-MCQs dataset, which has 113 multiple-choice questions from a first-year introductory materials science and engineering course at the University of Toronto. The questions are meant to test students’ understanding of materials science knowledge, including material mechanics, thermodynamics, crystal structures, materials properties, etc. The questions are categorized into easy (number of questions, n=39), medium (n=40), and hard levels (n=34), based on the conceptual difficulty, calculation complexity, and the depth of reasoning required. For the performance evaluation of property prediction, we use matbench_steels, a subcategory of the Matbench test set originally proposed for benchmarking traditional machine learning (ML) models for materials property predictions [21]. The matbench_steels dataset has 312 pairs of material compositions (as chemical formula) and yield strength (in MPa). For the robustness analysis of property prediction, we use a band gap dataset, which comprises 10,047 descriptions of material crystal structures generated via Robocrystallographer, along with their corresponding band gap values from the Materials Project database [14].
@@
Figure 1: Schematic representation of the experiment design for performance evaluation and robustness analysis of LLMs. Yellow highlights the testing conducted in Q&A settings. Blue highlights the testing conducted in property prediction settings. Green represents the tests associated with performance evaluation. Red represents tests related to degradation and robustness analysis.
Performance Evaluation of Materials Science Q&A
The results of the performance evaluation of LLMs on the MSE-MCQs dataset are shown in Figure 2. For each model and setting, three trials were conducted, and the error bars represent the inherent non-determinism in LLMs even at the lowest temperature settings. This non-determinism may affect the reliability and reproducibility of the models [22] and is likely a result of the stochastic sampling during text generation [23, 24, 25]. The result shows that gpt-4-0613 scored the highest in all categories, followed by gpt-3.5-turbo-0613 and llama2-70b. The remaining models scored around or slightly below the baseline score of 0.246, which is equivalent to random guessing for the provided MCQs.
Upon implementing the expert prompt, we observed a consistent improvement in model performance across almost all the models and question types. For gpt-3.5-turbo-0613, the significant variation in the data indicated by the large error bar prevents a definitive conclusion about the expert prompt improving performance on easy and medium-level questions. However, the accuracy of hard-level questions nearly doubled.
We investigated why the smaller Llama models (llama-13b and llama-7b) scored lower than the baseline without the expert prompt. Beyond the limited skill levels, these models even failed to understand the intent when instructions were not provided. Instead of answering, they often attempted to “complete” the questions rather than answering them. However, once the expert prompt was implemented, these smaller models could follow the instructions and attempt to solve the questions, in which case the performance improved to around and sometimes above the baseline scores.
The performance trends observed when evaluating large language models on the MSE-MCQs dataset were expected. Specifically, larger models like gpt-4-0613 significantly outperformed smaller ones, with model performance generally improving across the board when an expert prompt was used, especially in handling more complex questions.
Performance Evaluation of Materials Property Prediction
Here we investigate LLM materials property predictions, utilizing the matbench_steels dataset. Figure 3 presents a comparative analysis between gpt-3.5-turbo-0613, employing few-shot ICL, and a traditional random forest regressor (RFR) model, in predicting the yield strength of steels. It is important to note that gpt-3.5-turbo-0613 was not subject to an explicit training phase for this task but processed the data as few-shot examples. In contrast, the RFR was trained explicitly using MAGPIE features from the same data points [26].
Figure 3: Prediction performance vs. number of training points of RFR and gpt-3.5-turbo-0613. The distributions indicate models’ sensitivity to the selection of training points, both existing in the LLM model with ICL and RFR model with traditional training process. In each violin, the center dashed line represents the median, and the upper and lower dashed lines represent the first and third quartile, respectively.
Figure 3 summarizes the distribution of errors for the two models as a function of the number of training points used. It can be observed that gpt-3.5-turbo-0613, utilizing its ICL capabilities, performs comparably to the trained RFR model. When given only 5 data points as few-shot examples, the median MAE of gpt-3.5-turbo-0613 begins at a slightly higher value of 249 MPa but improves significantly to 171 MPa when increased to 25 data points. Meanwhile, the RFR model, starting with a slightly lower median MAE of 231 MPa with a large error bar, shows a more modest improvement to 194 MPa at 25 training points. The declining MAE trend suggests that both models benefit from more data, yet the LLM’s rate of improvement is more significant.
This analysis reveals a notable adaptability of pretrained LLMs to new predictive challenges using ICL, particularly when data availability is constrained. Such adaptability makes LLMs potentially valuable in early-stage research or exploratory studies in materials science, where data may be scarce or costly to obtain. However, as the number of data points increases, most LLMs suffer from limited prompt windows, which make such applications computationally expensive or impossible, in which case fine-tuned LLMs and traditional machine learning models with dedicated training may be more effective.
To better understand how the selection of highly relevant training data can be used to enhance LLM performance we conducted a systematic study using nearest-neighbor-boosted ICL. The method developed here involves deliberately selecting training data points based on representational proximity to enhance the model’s performance. Figure 4 shows the prediction performance when LLMs are tasked with predicting yield strength using a) farthest neighbors, b) random neighbors, and c) nearest neighbors. For performance comparison, a k-nearest neighbors (KNN) model and RFR were also implemented.
Figure 4: Prediction performance of various LLMs under different training neighbor settings. Three panels compare the model performance across neighbor selection methods (left to right): Farthest Neighbors, Random Neighbors, and Nearest Neighbors. The TPOT-Mat performance is indicated by a horizontal grey line for benchmarking.
When trained with farthest neighbors, models generally exhibit a decrease in MAE as the number of neighbors increases. A notable exception to this trend is llama2-13b, which shows a non-monotonic decrease in MAE. The GPT models both outperformed the KNN and RFR models when fed with more than 15 points. This indicates that LLMs may have an inherent capability to extract valuable patterns even from data points that are less obviously related to the test set.
For the random neighbors training set, the LLMs’ performance consistently improves as the dataset size increases. This suggests that a randomly composed dataset offers a balanced and rich learning environment for these models, allowing for the development of more robust generalizations about the data. The GPT models outperform the RFR model as the data size increases, indicating that their more sophisticated architectures and larger training corpora enhance their ability to analyze and interpret complex data relationships more effectively. On the other hand, the smaller LLMs exhibit less stable performance with higher MAE values which exhibit non-monotonic variation with the number of neighbors. The random neighbors setting appears to challenge these models to a greater degree, potentially because their smaller scale and limited pretraining are less capable of handling the diverse input without fine-tuning or additional data processing.
The nearest neighbors represent the most relevant data points in the compositional space to the test points. The KNN model exhibits an increase in MAE with a growing number of neighbors, which is expected as additional (more distant) neighbors are less similar to the predicted sample. This suggests that if LLMs act purely on information provided without additional internal calculation, their performance should degrade as the number of neighbors increases. The Llama models do not perform as well as the KNN or RFR models. Contrastingly, as the data size increases, the larger LLMs show a consistent decrease in MAE and outperform the RFR model. While the gpt-4-0613 model with 20 nearest neighbors achieves an MAE of 89.6, it does not surpass the best-performing ML model, TPOT-Mat, which has an MAE of 79.9 on the matbench_steels dataset [27]. However, it is important to note that this is not a direct comparison, as TPOT-Mat employs a 5-fold nested cross-validation method on Matbench datasets [21], which utilizes all available data and is likely to result in better performance. This highlights the promise of LLMs in data-lean materials property prediction tasks without the need for feature engineering, particularly given their general-purpose design and lack of task-specific fine-tuning.
To investigate the model’s predictive behaviors under these different settings, we analyzed the parity plots of the gpt-4-0613’s predictions when utilizing 20 neighboring data points, as shown in Figure 5. Alongside these plots, the figure also includes histograms of the top five most frequently predicted yield strength values, to ensure the model is not merely guessing a few commonly present values. This behavior is known as “mode collapse”, whereby a generative model can favor a certain output due to overfitting to its pretraining data or lack of generalization capability [28]. Understanding mode collapse is crucial for evaluating the robustness of LLMs because it directly impacts the model’s reliability and utility in practical applications. By identifying the mode collapse behavior, one can evaluate the quality of those predictions and potentially improve the performance.
Figure 5: Parity plots (top) and associated mode collapse point histograms (bottom) of gpt-4-0613 under different training neighbor settings: Farthest Neighbors, Random Neighbors, and Nearest Neighbors, from left to right.
In the farthest neighbors setting, the discrete nature of the predictions, as evidenced by the observation of horizontal lines in the figure, suggests that the model is unable to capture the relationship between the composition and the yield strength effectively. The histogram indicates a strong mode collapse behavior, with the model frequently predicting a handful of values. This suggests that the model may be defaulting to a “safe” prediction range (possibly from the pretraining data) when provided with less relevant examples. In the random neighbors setting, the model shows better overall performance and a reduced mode collapse behavior. This suggests that introducing more variability into the training data helps the model to better understand the underlying patterns that predict yield strength. The nearest neighbors setting exhibits the best model performance suggesting that proximity in training data correlates with more accurate yield strength predictions. The mode collapse behavior is significantly reduced compared to the farthest neighbors and random neighbors, reflecting a greater diversity in the model’s output. The observations show varying degrees of mode collapse based on the proximity of prompt examples to the predicted point. For instance, when provided with more closely related training data, the model exhibits stronger predictive signals. This suggests that the extent of mode collapse behavior could serve as an indicator of LLMs’ predictive quality, which can be improved by including more similar examples in the prompt.
Robustness Analysis of Materials Science Q&A
In Figure 6, we present the outcomes of the robustness assessment of gpt-3.5-turbo-0613 when confronted with different types of textual modifications to the MSE-MCQs, to evaluate its stability to various types of "noise". The gpt-3.5-turbo-0613 was chosen because of its moderate performance, availability, and cost-effectiveness.
Figure 6: Robustness analysis of gpt-3.5-turbo-0613 in materials science Q&A using MSE-MCQs dataset. The error bars represent the standard deviation due to LLMs’ non-determinism (Original, Synonym Replacement, Distractive Info) and the randomness (Sentence Reordering, Unit Mixing, Superfluous Info) introduced to the questions.
Ranking by the degradation severity on the easy-level questions, sentence reordering has the least performance drop, followed by synonym replacement, distractive info, unit mixing, and superfluous info. The performance on the hard-level questions is close to the baseline score, indicating the model’s struggle with complex queries regardless of textual modifications, and will not be discussed in detail. The larger error bars in medium and hard questions suggest that LLMs tend to generate more varied responses to complex and lengthy queries.
Sentence reordering has a minimal impact on gpt-3.5-turbo-0613’s performance, demonstrating the model’s robust parsing capabilities. The model maintains comprehension and extracts the necessary information to answer correctly, despite the alteration in the natural flow of questions. This suggests an advanced level of syntactic flexibility and the ability to contextualize information effectively.
The performance drop with synonym replacement indicates that the LLMs can generate different or faulty responses when presented with different terminologies of the same meaning. This reveals LLMs’ reliance on specific terminologies for recognition and comprehension in materials science. In contrast to humans, who can grasp the conceptual continuity behind varied expressions for flexible cognition, LLMs’ struggles with synonym replacement emphasize the need for advanced training that prioritizes semantic networks over mere word recognition [29].
Introducing distractive information mimics a real-world scenario where irrelevant data often accompanies critical information, requiring sharp focus and analytical precision. Improving LLMs’ ability to filter out irrelevant information is crucial for more effective information retrieval, problem-solving, and data interpretation [30]. The results indicate that LLMs struggle to filter distractions in easy-level questions, where mostly simple conceptual knowledge is tested. The added non-essential details seemed to divert the LLMs’ "attention", leading to fewer correct answers. The degradation is less severe in the medium and hard-level questions, where greater complexities and lengths may make the distractive information more distinguishable.
Mixing and converting the units tests LLMs’ abilities to perform numerical reasoning and apply mathematical concepts within a linguistic context. The added complexity introduced by unit mixing degraded the performance of the gpt-3.5-turbo-0613 model on easy and medium-level questions. Although some state-of-the-art LLMs support multi-modal applications and function calls to perform calculations [31], the ability to correctly identify and convert units from a large text can still benefit use cases such as information retrieval and data interpretation.
Superfluous information differs from distractive information in that it is more relevant to the questions themselves. For LLMs, distinguishing the necessary information from merely relevant but non-essential details is a more challenging cognitive process, mirroring the nuances of advanced human problem-solving. It requires an understanding of the problem’s objective, prioritizing information based on the question, and applying only the information that will lead to the correct conclusion. The noted decrease in accuracy shows a current limitation in the LLM’s ability for critical information assessment. The variation in performance degradation is influenced by the type and relevancy of the superfluous information provided.
Robustness Analysis of Materials Property Prediction
Original Distractive Info Sentence Reordering Superfluous Info
[…]
Table 1: Mean Absolute Error (MAE) under different conditions
The impact of sentence reordering increased the MAE by 12.9%, suggesting the model’s reliance on the structured descriptions for accurate predictions. From the previous study on MSE-MCQs degradation, the effect of sentence reordering was less significant, indicating that larger general LLMs, which are trained with more various texts, can exhibit better contextual understanding and are less prone to order changes. The presence of superfluous information, particularly an additional sentence from another material’s description, leads to a 39% increase in MAE. This substantial degradation indicates that while the model can filter out irrelevant distractive noise, it struggles considerably when faced with data that is contextually relevant to the specific prediction task.
To further assess the model’s robustness and determine which description elements are essential for prediction accuracy, we conducted an ablation study that involves altering the structures and lengths of the input description. As shown in Figure 7, the description length is expressed as percent sentence inclusion, ranging from 10% to 100% and MAE is used as a measure of prediction accuracy.
Figure 7: Structure and length effect of LLM-Prop on prediction performance.
When the number of description sentences is incrementally increased, the MAE rapidly decreases as more information is provided, and optimized at 100% sentence inclusion. Interestingly, in the other configurations, the initial MAEs with just 10% sentence inclusion are notably lower and some can double the performance of the original setting. This indicates that the initial sentences may not contain the most useful information for prediction. The MAEs begin to converge around 50% sentence inclusion, beyond which differences become statistically insignificant. Notably, in the random order setting, there is virtually no variation in the MAE when incorporating three different sentence shuffles. This suggests that LLM-Prop can effectively extract key information and deliver consistent predictions, despite variations in the sentence order.
By 40% sentence inclusion, the reversed order yields the lowest MAE, indicating that sentences at the end of descriptions contain crucial information. However, by 50% sentence inclusion, the performance of the reversed order begins to align with that of the random order. This alignment suggests that central information may not be as crucial for prediction accuracy, especially as the random order likely incorporates more of the initial sentences than the reversed order at this inclusion level. Based on these insights, we developed the sides-to-middle approach, aiming to prioritize information at the beginning and the end. The approach shows the most significant improvement, achieving the lowest MAE between 40% and 70% sentence inclusion among all. The error continues to decrease and is optimized at full sentence inclusion being only 5.8% higher than the original setting in MAE. This implies that the contextual framing provided by the beginning and end of the description holds considerable importance for the model’s predictive accuracy, yet the model is still best optimized for the original order.
This ablation study showcases that the fine-tuned model can perform effectively even when provided with significantly reduced prompts. We found that diverging from the training setup (i.e., changing the textual order of the prompt) can sometimes result in improved performance at ablated data volumes. This counterintuitive result suggests that highly templated training or fine-tuning data can lead to unexpected effects. Consequently, this implies two key considerations: 1) training templates should be diverse to prevent models from overfitting to unimportant patterns, and 2) when using a fine-tuned model trained on a specific template, it may not always be optimal to match the template during inference, particularly under ablated data conditions. These insights highlight the potential for optimizing training costs while maintaining performance.
Discussion
The findings of this study offer crucial insights into the behaviors and limitations of LLMs in materials science domain-specific Q&A and materials property prediction. While the robustness analysis indicates that LLMs can manage certain types of noise with resilience, their performance is significantly challenged under more complex and deceptive conditions, such as when superfluous information is introduced. Prompting techniques (e.g. expert prompting, zero-shot chain-of-thought prompting) can sometimes improve the model performance, not by unlocking new capabilities, but by increasing the probability of following an expected format. During property prediction, in-context learning enables pretrained LLMs to achieve relatively high accuracy in low-data settings when provided with few-shot examples with high proximity to the target. However, the observed mode collapse behavior, where the model generates repeated outputs despite varying inputs, showcases that providing ineffective few-shot examples can cause the model to default to a memorized response rather than conditioning its output on the provided prompt. This study also highlights that fine-tuned models can exhibit enhanced performance under ablated data conditions when diverging from the training setup, such as altering the textual order. This underscores the risks associated with fixed templating during fine-tuning and suggests that users should be cautious about strictly matching the training templates during inference. This study highlights the challenges and limitations of using LLMs in materials science, aiming to motivate future improvements that enhance their reliability and effectiveness as research tools.
Methods
The methodology is divided into four subsections that cover the performance evaluation and robustness analysis of LLMs in materials science Q&A and property predictions. In each subsection, we will introduce the models, datasets used, prompting techniques, and evaluation criteria chosen for the specific study.
Performance Evaluation of Materials Science Q&A

### v2

@@ Abstract
Large Language Models (LLMs) have the potential to revolutionize scientific research, yet their robustness and reliability in domain-specific applications remain insufficiently explored. In this study, we evaluate the performance and robustness of LLMs for materials science, focusing on domain-specific question answering and materials property prediction across diverse real-world and adversarial conditions. Three distinct datasets are used in this study: 1) a set of multiple-choice questions from undergraduate-level materials science courses, 2) a dataset including various steel compositions and yield strengths, and 3) a band gap dataset, containing textual descriptions of material crystal structures and band gap values. The performance of LLMs is assessed using various prompting strategies, including zero-shot chain-of-thought, expert prompting, and few-shot in-context learning. The robustness of these models is tested against various forms of ‘noise’, ranging from realistic disturbances to intentionally adversarial manipulations, to evaluate their resilience and reliability under real-world conditions. Additionally, the study showcases unique phenomena of LLMs during predictive tasks, such as mode collapse behavior when the proximity of prompt examples is altered and performance recovery from train/test mismatch. The findings aim to provide informed skepticism for the broad use of LLMs in materials science and to inspire advancements that enhance their robustness and reliability for practical applications.
1 Department of Materials Science and Engineering, University of Toronto, Canada.
2 Acceleration Consortium, University of Toronto, Canada.
3 Artificial, Inc., Austin, Texas, USA.
Corresponding authors: edwardsoo.kim@mail.utoronto.ca, jason.hattrick.simpers@utoronto.ca
1   Introduction
@@
Large Language Models (LLMs) represent a significant advancement in the field of artificial intelligence and have been rapidly adopted for application in various scientific disciplines [1]. With their ability to process and generate natural language, LLMs are potent tools for tasks like information retrieval, question and answering (Q&A), and property predictions [2, 3, 4]. Similar to traditional ML models, LLMs can require extensive data processing, large volumes of data, and massive compute resources to train [5]. Despite these limitations, pretrained LLMs can be adapted to new tasks with few-shot examples via in-context learning (ICL), making them both cost-effective and rapid to deploy [6, 7, 8]. In the context of materials science, where data acquisition can often be costly and time-consuming, leveraging ICL enables LLMs to efficiently prototype and generate predictive insights even in low-data settings [9, 10, 11]. Recent work has demonstrated that LLMs are capable of domain-specific Q&A [12, 13, 14], materials property predictions [15, 16, 17], and information extraction from complex datasets [4, 18]. These studies demonstrate LLMs’ potential to serve as flexible and powerful analytical tools.
@@
However, the robustness of LLMs is a critical factor in their practical deployment, yet it remains an underexplored area, particularly in domain-specific applications such as materials science. Previous studies have shown that LLMs struggle to maintain predictive accuracy when the input distribution shifts, exhibiting poor generalization to out-of-distribution (OOD) test data and vulnerability to adversarial attacks [19, 20, 21]. These challenges highlight the need for systematic robustness evaluations to ensure LLM reliability in real-world scenarios. A key aspect of the robustness of LLMs is their sensitivity to prompt changes either due to innocuous or adversarial reasons [22, 23]. Variations in how a query or instruction is formulated may cause a response to factually change [22]. As an example, 0.1 nm and 1 Angstrom are equivalent but switching them in a prompt could result in different LLM predictions for the same task. Alternatively, the response of the LLM can be deliberately altered through intentional misinformation or misleading inputs [23]. These attributes are not only theoretical concerns but are critical for the reliable usage of LLMs as they become integrated into the materials science research and development pipeline. Given that LLMs generate outputs with indifference to truth [24], thoroughly probing LLM prompt sensitivity would allow us to critically evaluate model performance in practical situations; providing informed skepticism for the broad use of LLMs in materials science.
@@
In this work, we conducted a holistic robustness analysis of commercial and open-source LLMs for materials science. While our primary analyses focus on pre-reasoning models due to their consistent single-pass inference structure, we also include a representative reasoning model (DeepSeek-R1[25]) in our initial benchmarking to contextualize performance differences. Reasoning models, such as DeepSeek-R1 and OpenAI-o1[26], dynamically adjust inference strategies based on problem complexity, introducing variability that complicates direct comparisons with single-pass models. To maintain consistency and isolate inherent robustness characteristics, we restrict our subsequent robustness and sensitivity analyses solely to pre-reasoning LLMs. Three distinct datasets of domain-specific Q&A and materials property prediction were selected. First, we benchmarked LLMs of different sizes and release periods using prompt engineering to establish baseline and optimal performance boundaries. We then investigated the impact of various textual perturbations, ranging from realistic to adversarial, on LLM performance in materials science Q&A. We then used the matbench_steels dataset to show that even without fine-tuning, pretrained LLMs demonstrated enhanced predictive ability through few-shot ICL when presented with similar examples to the prediction task. Conversely, when provided with dissimilar examples during few-shot ICL, mode collapse behavior was observed, where the model generated identical outputs despite varying inputs. Counterintuitively, for the fine-tuned LLM examined on band gap predictions, supposedly adversarial perturbations like shuffling or randomization enhanced the model’s predictive capability. Interestingly, this enables the fine-tuned model to function effectively even with significantly truncated prompts. This train/test mismatch behavior, absent in traditional ML models, highlights a potential direction for distilling LLM-based predictive models.
2   Methods
@@
The methodology is divided into four subsections that cover the performance evaluation and robustness analysis of LLMs in materials science Q&A and property predictions. In each subsection, we will introduce the models, datasets used, prompting techniques, and evaluation criteria chosen for the specific study. All the evaluated models were set to their lowest temperature (typically 0) to minimize the non-determinism and maximize reproducibility. Figure 1 illustrates the experimental framework for evaluating LLMs in materials science Q&A and materials property prediction. For performance evaluation and robustness analysis of materials science Q&A, we compiled the MSE-MCQs dataset, consisting of 113 multiple-choice questions specifically designed for this study. These questions are original and were created by faculty at the University of Toronto for a first-year introductory materials science and engineering course. The questions were designed to test students’ understanding of materials science knowledge, including material mechanics, thermodynamics, crystal structures, materials properties, etc. The questions are manually categorized into easy (number of questions, n=39), medium (n=40), and hard levels (n=34), based on a set of heuristics, including conceptual complexity, the level of reasoning required, and the presence and difficulty of the calculations. For example, "easy" questions primarily test factual recall or direct application of basic concepts, such as identifying the crystal structure of a material. "Medium" questions involve moderate reasoning or straightforward calculations, such as determining the stress in a material under specific conditions. "Hard" questions require multi-step reasoning or more complex calculations, such as deriving material properties from combined thermodynamic and mechanical data. For the performance evaluation of property prediction, we use matbench_steels, a subcategory of the Matbench test set originally proposed for benchmarking traditional machine learning (ML) models for materials property predictions [27]. The matbench_steels dataset has 312 pairs of material compositions (as chemical formula) and yield strength (in MPa). For the robustness analysis of property prediction, we use a band gap dataset, which comprises 10,047 descriptions of material crystal structures generated via Robocrystallographer, along with their corresponding band gap values from the Materials Project database [15].
@@
Figure 1: Schematic representation of the experiment design for performance evaluation and robustness analysis of various LLMs. Yellow highlights the testing conducted in Q&A settings. Blue highlights the testing conducted in property prediction settings. Green represents the tests associated with performance evaluation. Red represents tests related to degradation and robustness analysis.
2.1   Performance Evaluation of Materials Science Q&A